OpenStudy (rayford):

identify the discontinuities, graph, and find the zeros of each rational function f(x)=3x^2-7x-6/x-3

OpenStudy (anonymous):

hint: x-3=/0 =>x=/ 3 3x^2-7x-6=0

OpenStudy (rayford):

wait come back!! i suck horribly AT THIS!

OpenStudy (anonymous):

u need to factor that now

OpenStudy (rayford):

ok..so...

OpenStudy (rayford):

i have been stuck with this for hours.... could you just work it out and let me review?

OpenStudy (anonymous):

alright hold on

OpenStudy (anonymous):

alright here's what i did: i used Delta for factoring ,delta= b^2-4(c)(a) then for x: x= -b+or- sqr of delta/2a i got x=3 or x=-2/3

OpenStudy (anonymous):

the delta formula is used to solve equations from lvl 2,which means that have x^2

OpenStudy (anonymous):

u can use that for factoring and u get (x-3)(x+2/3)